Jim's nephew owns a triangular plot of land. The perimeter fence that surounds the land measures 121 feet. The shortest side is 15 feet longer than one half of the longest side. The second longest side is 4 feet shorter than the longest side. What is the length of each side.
Answer by MathTherapy(10858)

Let the longest side = l
Then shortest side = 15 + .5l
And, second longest side = l - 4
Since the perimeter is 121,” then l + 15 + .5l + l - 4 = 121
2.5l = 110
l = 44
Therefore, longest side = 44 feet
Second longest = 44 - 4 = 40 feet
Shortest side = 15 + .5(44) = 37 feet